# Observing structural disorder induced interacting topological phase in an atom array

### v2

@@ Acknowledgements.
In the following, we will show how the energy shift at the edge sites caused by van der Waals interactions lift the four-fold degeneracy. Let $|\Psi(N-1)\rangle$ and $|\Psi(N+1)\rangle$ be the ground states of $H_{\mathrm{XXZ}}$ in the subspace with $N-1$ and $N+1$ particles, respectively, and $|\Psi_{L}(N)\rangle$ and $|\Psi_{R}(N)\rangle$ be the two ground states of $H_{\mathrm{XXZ}}$ in the half-filling subspace. For $H_{\mathrm{XXZ}}$, when it is topologically nontrivial, the four ground states are degenerate. However, the $H_{\mathrm{Z}}$ term modifies the spectra. For simplicity, we consider its first-order correction for a ground state $|\Psi\rangle$, i.e., $\Delta E_{\Psi}=\langle\Psi|H_{\mathrm{Z}}|\Psi\rangle=-\frac{1}{4}\sum_{i=1}^{2N}V_{i}^{\text{vdW}}(2\langle\Psi|\hat{n}_{i}|\Psi\rangle-1)$, where we have expressed $H_{\mathrm{Z}}=-\frac{1}{4}\sum_{i=1}^{2N}V_{i}^{\text{vdW}}(2\hat{n}_{i}-1)$ in terms of $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$, the particle number operator at site $i$. For the four states involved, in the bulk sites with $12\leq i\leq 2N-$ 1, we approximate $\langle\Psi|\hat{n}_{i}|\Psi\rangle\approx 0.5$ so that $\Delta E_{\Psi}\approx-\frac{1}{4}\left[V_{1}^{\text{vdW}}(2\hat{n}_{1}-1)+V_{2N}^{\text{vdW}}(2\hat{n}_{2N}-1)\right]$. For $|\Psi_{L}(N)\rangle$ and $|\Psi_{R}(N)\rangle$, there is a particle occupying either the left or the right edge, respectively. We thus have $\langle\Psi_{L}(N)|\hat{n}_{1}|\Psi_{L}(N)\rangle\approx 1$, $\langle\Psi_{L}(N)|\hat{n}_{2N}|\Psi_{L}(N)\rangle\approx 0$, $\langle\Psi_{R}(N)|\hat{n}_{1}|\Psi_{R}(N)\rangle\approx 0$, and $\langle\Psi_{R}(N)|\hat{n}_{2N}|\Psi_{R}(N)\rangle\approx 1$. As a result, $\Delta E_{\Psi_{L}(N)}\approx-\frac{1}{4}(V_{1}^{\text{vdW}}-V_{2N}^{\text{vdW}})$ and $\Delta E_{\Psi_{R}(N)}=-\Delta E_{\Psi_{L}(N)}$. We see that for a typical realization, $\Delta E_{\Psi_{L}(N)}\neq\Delta E_{\Psi_{R}(N)}$, implying that the degeneracy is lifted. However, if we consider the inversion partner, then $\Delta E_{\Psi_{L}(N)}$ becomes $\frac{1}{4}(V_{1}^{\text{vdW}}-V_{2N}^{\text{vdW}})$ and $\Delta E_{\Psi_{R}(N)}$ becomes $-\frac{1}{4}(V_{1}^{\text{vdW}}-V_{2N}^{\text{vdW}})$. The average local DOS at the two edges remains equal at the same energy, which we have proven in the previous subsection to show that the two states are statistically degenerate. For $|\Psi(N-1)\rangle$, since no particles occupy the edges, $\langle\Psi(N-1)|\hat{n}_{1}|\Psi(N-1)\rangle\approx 0$ and $\langle\Psi(N-1)|\hat{n}_{2N}|\Psi(N-1)\rangle\approx 0$. Thus, $\Delta E_{\Psi(N-1)}=\frac{1}{4}(V_{1}^{\text{vdW}}+V_{2N}^{\text{vdW}})$. For $|\Psi(N+1)\rangle$, there are two particles occupying the edges so that $\langle\Psi(N+1)|\hat{n}_{1}|\Psi(N+1)\rangle\approx 1$ and $\langle\Psi(N+1)|\hat{n}_{2N}|\Psi(N+1)\rangle\approx 1$. Thus, $\Delta E_{\Psi(N+1)}=-\frac{1}{4}(V_{1}^{\text{vdW}}+V_{2N}^{\text{vdW}})$. We therefore conclude that van der Waals interactions lift the degeneracy between the two ground states $\Delta E_{\Psi(N-1)}$ and $\Delta E_{\Psi(N+1)}$, with their energy difference of about $\frac{1}{2}|V_{1}^{\text{vdW}}+V_{2N}^{\text{vdW}}|$.
@@
Figure S1(D) shows that $V_{1}^{\text{vdW}}/h,V_{2N}{{}^{\text{vdW}}}/h\sim 0.04$ MHz, suggesting an energy difference of about $0.04$ MHz. However, our numerical results show that the two ground states exhibit an energy difference of about $0.2$ MHz on average. This can be probably attributed to the distribution of edge states inside the bulk sites, where energy shifts by van der Waals interactions are much larger than at the edges. The energy difference between the ground state in the subspace with $N-1$ and $N$ particles is about $0.1$ MHz, which is consistent with our experimental observation that an occupancy peak at two edges arises at the detuning of $-0.1$ MHz (see Fig. 4A in the main text).
III.2.7 S3.2.7. Adiabatic preparation of the many-body ground state
Here we present numerical analyses of the adiabatic preparation for the many-body ground state in a subspace with a fixed particle number. As stated in the main text, we first prepare a product state $\left|\psi_{0}\right\rangle$, such as $|1010...10\rangle$, $|0010...10\rangle$, or $|1010...11\rangle$, which lie in the subspaces of $N$ (half-filling), $N-1$, and $N+1$ particles, respectively. Subsequently, we evolve it by the following Hamiltonian,
$\hat{H}_{\mathrm{tot}}=\hat{H}-\Delta E(t)\sum_{i\in S}\hat{n}_{i},$ (S28)
where $S$ represents the set of sites with applied pinning lasers. Since $\Delta E(t=0)/h=20$ MHz is much larger than the energy scale in $\hat{H}$, the state $\left|\psi_{0}\right\rangle$ is a good approximation to the ground state of $\hat{H}_{\mathrm{tot}}$. During the time evolution, we slowly ramp down the amplitude of the pinning lasers so that $\Delta E(t)=\Delta Ee^{-t/\tau}$ with $\tau=0.6\text{ }\mathrm{\mu s}$. The total evolution time is 4 $4\text{ }\mathrm{\mu s}$.
The effectiveness of the adiabatic evolution depends on the low energy spectral gap. Figure S2 displays this gap during the time evolution. We see that for the regular configuration, the gap quickly decreases, reaching a minimum value of $\Delta_{\mathrm{g}}/h=1.06$ MHz. For disordered configurations, the minimum energy gaps averaged over random configurations are $0.75$ MHz, $0.89$ MHz, and $0.72$ MHz for $N$, $N-1$, and $N+1$ particles, respectively, which are slightly smaller than the minimum energy gap in the regular case. The energy gaps are on the order of the energy scale of the hopping term in the Hamiltonian $\hat{H}$. At half-filling, it is defined as the gap between the third lowest energy state with respect to the ground state, as the ground state is nearly two-fold degenerate. During the time evolution, although the gap declines, the ramping speed for the pinning laser beams decreases correspondingly, thereby ensuring a successful adiabatic evolution.
Figure S2 also plots the fidelity between the evolving state and the corresponding true ground state during the time evolution without accounting for errors, where the evolving state is obtained by simulating the time evolution of the Hamiltonian using the Krylov method. We see that at the end of the adiabatic evolution, the fidelity reaches $0.99$ for the regular configuration. For the disordered cases, their average fidelity values are $0.89$, $0.75$, and $0.72$ for the fillings of $N$, $N-1$, and $N+1$ particles, respectively. To ensure that imperfections of the experiments are well understood, we also perform simulations by considering all SPAM errors. The results of numerical simulations are found to be in good agreement with the experimental data, as demonstrated in Fig. 3 in the main text.
@@
Figure S3: The residual edge spin magnetization for the quench dynamics in the regular configuration. (A) Level-spacing statistics for the Hamiltonian in a regular lattice at half-filling with inversion symmetry resolved. The expected distributions for an integrable model (Poisson) and the corresponding chaotic or thermalized model (Wigner-Dyson) are plotted. The energy level spacing $s$ is computed using the middle 20% of all eigenstates and rescaled such to $\langle s\rangle=1$. (B) Magnetization from numerical simulation of quench dynamics of the left edge spin with and without the $H_{\mathrm{Z}}$ term. The dashed line represents the predicted value in the diagonal ensemble and the right edge spin shows similar long time behavior.
III.2.8 S3.2.8. Effects of van der Waals interactions on quench dynamics
@@
Without the $H_{\mathrm{Z}}$ term, we observe that the initially polarized local spins delocalize and spread across the chain due to spin exchange processes, i.e., a particle can hop to other sites in the hard-core boson language. Over time, the spin distribution becomes homogeneous, leading to an average magnetization of $\langle\sigma_{i}^{z}\rangle=0$ at every site (see Fig. S3B). In the presence of a homogeneous $H_{\mathrm{Z}}$ term, the physics remains unchanged since the dynamics are restricted to the subspace of $\sigma^{z}=0$. This indicates that the slight residual arises from the difference in the $H_{\mathrm{Z}}$ term between the edges and the bulk. Due to this difference in onsite potential, bosons at the edges prefer to move into the bulk, with the energy decrease compensated for by Ising interactions between the particles.

### v3

@@ Acknowledgements.
In the following, we will show how the energy shifts at the edge sites caused by van der Waals interactions lift the four-fold degeneracy. Let $|\Psi(N-1)\rangle$ and $|\Psi(N+1)\rangle$ be the ground states of $H_{\mathrm{XXZ}}$ in the subspace with N1N-1 and $N+1$ particles, respectively, and $|\Psi_{L}(N)\rangle$ and $|\Psi_{R}(N)\rangle$ be the two ground states of $H_{\mathrm{XXZ}}$ in the half-filling subspace. For $H_{\mathrm{XXZ}}$, when it is topologically nontrivial, the four ground states are degenerate. However, the $H_{\mathrm{Z}}$ term modifies the spectra. For simplicity, we consider its first-order correction for a ground state $|\Psi\rangle$, i.e., $\Delta E_{\Psi}=\langle\Psi|H_{\mathrm{Z}}|\Psi\rangle=-\frac{1}{4}\sum_{i=1}^{2N}V_{i}^{\text{vdW}}(2\langle\Psi|\hat{n}_{i}|\Psi\rangle-1)$, where we have expressed $H_{\mathrm{Z}}=-\frac{1}{4}\sum_{i=1}^{2N}V_{i}^{\text{vdW}}(2\hat{n}_{i}-1)$ in terms of $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$, the particle number operator at site $i$. For the four states involved, in the bulk sites with $2\leq i\leq 2N-1$, we approximate $\langle\Psi|\hat{n}_{i}|\Psi\rangle\approx 0.5$ so that $\Delta E_{\Psi}\approx-\frac{1}{4}\left[V_{1}^{\text{vdW}}(2\hat{n}_{1}-1)+V_{2N}^{\text{vdW}}(2\hat{n}_{2N}-1)\right]$. For $|\Psi_{L}(N)\rangle$ and $|\Psi_{R}(N)\rangle$, there is a particle occupying either the left or the right edge, respectively. We thus have $\langle\Psi_{L}(N)|\hat{n}_{1}|\Psi_{L}(N)\rangle\approx 1$, $\langle\Psi_{L}(N)|\hat{n}_{2N}|\Psi_{L}(N)\rangle\approx 0$, $\langle\Psi_{R}(N)|\hat{n}_{1}|\Psi_{R}(N)\rangle\approx 0$, and $\langle\Psi_{R}(N)|\hat{n}_{2N}|\Psi_{R}(N)\rangle\approx 1$. As a result, $\Delta E_{\Psi_{L}(N)}\approx-\frac{1}{4}(V_{1}^{\text{vdW}}-V_{2N}^{\text{vdW}})$ and $\Delta E_{\Psi_{R}(N)}=-\Delta E_{\Psi_{L}(N)}$. We see that for a typical realization, $\Delta E_{\Psi_{L}(N)}\neq\Delta E_{\Psi_{R}(N)}$, implying that the degeneracy is lifted. However, if we consider the inversion partner, then $\Delta E_{\Psi_{L}(N)}$ becomes $\frac{1}{4}(V_{1}^{\text{vdW}}-V_{2N}^{\text{vdW}})$ and $\Delta E_{\Psi_{R}(N)}$ becomes $\frac{1}{4}(V_{1}^{\text{vdW}}-V_{2N}^{\text{vdW}})$. The average local DOS at the two edges remains equal at the same energy, which we have proven in the previous subsubsection to show that the two states are statistically degenerate. For $|\Psi(N-1)\rangle$, since no particles occupy the edges, $\langle\Psi(N-1)|\hat{n}_{1}|\Psi(N-1)\rangle\approx 0$ and $\langle\Psi(N-1)|\hat{n}_{2N}|\Psi(N-1)\rangle\approx 0$. Thus, $\Delta E_{\Psi(N-1)}=\frac{1}{4}(V_{1}^{\text{vdW}}+V_{2N}^{\text{vdW}})$. For $|\Psi(N+1)\rangle$, there are two particles occupying the edges so that $\langle\Psi(N+1)|\hat{n}_{1}|\Psi(N+1)\rangle\approx 1$ and $\langle\Psi(N+1)|\hat{n}_{2N}|\Psi(N+1)\rangle\approx 1$. Thus, $\Delta E_{\Psi(N+1)}=-\frac{1}{4}(V_{1}^{\text{vdW}}+V_{2N}^{\text{vdW}})$. We therefore conclude that van der Waals interactions lift the degeneracy between the two ground states $\Delta E_{\Psi(N-1)}$ and $\Delta E_{\Psi(N+1)}$, with their energy difference of about $\frac{1}{2}|V_{1}^{\text{vdW}}+V_{2N}^{\text{vdW}}|$.
@@
Figure S2d shows that $V_{1}^{\text{vdW}}/h,V_{2N}^{\text{vdW}}/h\sim 0.04$ MHz, suggesting an energy difference of about $0.04$ MHz. However, our numerical results show that the two ground states exhibit an energy difference of about $0.2$ MHz on average. This can be probably attributed to the distribution of edge states inside the bulk sites, where energy shifts by van der Waals interactions are much larger than at the edges. The energy difference between the ground states in the subspace with N1N-1 and $N$ particles is about $0.1$ MHz, which is consistent with our experimental observation that an occupancy peak at two edges arises at the detuning of 0.1-0.1 MHz (see Fig. 4a in the main text).
@@
Figure S3: The residual edge spin magnetization for the quench dynamics in the regular configuration. a, Level-spacing statistics for the Hamiltonian in a regular lattice at half-filling with inversion symmetry resolved. The expected distributions for an integrable model (Poisson) and the corresponding chaotic or thermalized model (Wigner-Dyson) are plotted. The energy level spacing $s$ is computed using the middle 20% of all eigenstates and rescaled such to $\langle s\rangle=1$. b, Magnetization from numerical simulation of quench dynamics for the left edge spin with and without the $H_{\mathrm{Z}}$ term. The dashed line represents the predicted value in the diagonal ensemble and the right edge spin shows similar long time behavior.
III.2.7 S3.2.7. Effects of van der Waals interactions on quench dynamics
@@
Without the $H_{\mathrm{Z}}$ term, we observe that the initially polarized local spins delocalize and spread across the chain due to spin exchange processes, i.e., a particle can hop to other sites in the hard-core boson language. Over time, the spin distribution becomes homogeneous, leading to an average magnetization of $\langle\sigma_{i}^{z}\rangle=0$ at every site (see Fig. S3b). In the presence of a homogeneous $H_{\mathrm{Z}}$ term, the physics remains unchanged since the dynamics are restricted to the subspace of $\sigma^{z}=0$. This indicates that the slight residual arises from the difference in the $H_{\mathrm{Z}}$ term between the edges and the bulk. Due to this difference in onsite potentials, bosons at the edges prefer to move into the bulk, with the energy decrease compensated for by Ising interactions between the particles.